Use the following contingency table:

A B

Yes 40 25

No 35 45

The calculated Chi-squared test statistic is

1. 5:45

2. 4:55

3. 145

4. 3:64

5. 45:45


1
Expert's answer
2021-08-16T15:29:38-0400


The chi-square test:

"χ^2 = \sum \frac{(O_i -E_i)^2}{E_i} \\

= \frac{(40-33.62)^2}{33.62} + \frac{(25-31.38)^2}{31.38} + \frac{(35-41.38)^2}{41.38} + \frac{(45-38.62)^2}{38.62} \\

= 1.21 + 1.30 + 0.98 + 1.05 \\

= 4.54"

Answer: option 2. 4:55